# Y. Grandati1 and C. Quesne2

### v1

@@ VII CONCLUSION
In this article, we have studied the multi-step version of a previous construction of the TDPT potential regular rational extensions obtained by one-step DT using seed functions associated with the para-Jacobi polynomials of Calogero and Yi. We have shown that the eigenstates of such $m$-step extensions are expressed in terms of novel families of EOPs, orthogonal on ]1,+1[]-1,+1[] - 1 , + 1 [ and depending not only on $m$ discrete parameters, but also on $m$ real continuous parameters $\lambda_{1}$, $\lambda_{2}$, …, $\lambda_{m}$. The sets of parameters are related by some restrictions coming from the TDPT rational extensions regularity conditions, which we study in detail.

### v2

@@ VII CONCLUSION
In this article, we have studied the multi-step version of a previous construction of the TDPT potential regular rational extensions obtained by one-step DT using seed functions associated with the para-Jacobi polynomials of Calogero and Yi. We have shown that the eigenstates of such $m$-step extensions are expressed in terms of novel families of EOPs, orthogonal on ]1,+1[]-1,+1[ and depending not only on $m$ discrete parameters, but also on $m$ real continuous parameters $\lambda_{1}$, $\lambda_{2}$, …, $\lambda_{m}$. The sets of parameters are related by some restrictions coming from the TDPT rational extensions regularity conditions, which we study in detail.